OpenStudy (anonymous):

Please help! Let v=(5,-2) and w=(-10,4). Which of the following are true? Check all that apply. a. w=-2v b. the y-component of w is 4. c. v*w=-58 d. v and w are perpendicular

OpenStudy (jdoe0001):

x y v = ( 5, -2) w = (-10, 4).

OpenStudy (anonymous):

okay but I still don't understand

OpenStudy (jdoe0001):

ok... well, what.... part? :)

OpenStudy (anonymous):

how to find the answer! ahah

OpenStudy (anonymous):

I think A is true?

OpenStudy (jdoe0001):

what's the "y" in the vector w?

OpenStudy (anonymous):

4

OpenStudy (jdoe0001):

b. the y-component of w is 4.

OpenStudy (anonymous):

oh okay! So A and B are true, how about the other ones? How do I multiply v and w together?

OpenStudy (jdoe0001):

now, -2<5, -2> => <(-2*5), (-2*-2)> => <-10, 4> = w

OpenStudy (jdoe0001):

yeah, the multiplication, so-called the dot product \(\bf <a, b> \cdot <c, d> \implies a\times c + b \times d\)

OpenStudy (anonymous):

so 5*-10+-2*4=-58

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so C is also true, do you know if they are perpendicular?

OpenStudy (jdoe0001):

2 vectors are perpendicular, or so-called orthogonal, if their dot product equals 0

OpenStudy (anonymous):

Okay so they are not! Thanks so, so much!! :)

OpenStudy (jdoe0001):

yw